OpenStudy (anonymous):

Remi has two cone-shaped containers with the same diameter. He will place the smaller container inside the larger one. Before he does this, he wants to fill the larger container with water so that it will be completely full but won’t spill when he places the smaller container inside. The diameter of both containers is 12 cm. The height of the smaller container is 6 cm, and the height of the larger container is 18 cm. What volume of water must Remi put in the large container? Use 3.14 to approximate pi and express your final answer in hundredths.

OpenStudy (anonymous):

@khalilforthewin

OpenStudy (khalilforthewin):

2034.72

OpenStudy (anonymous):

Isn't it 718.01?

OpenStudy (khalilforthewin):

where did you get that from

OpenStudy (anonymous):

It's ___cm3. I can't retake this again haha

OpenStudy (anonymous):

use the formula for the cone 1/3 times (area of base) times height so first we want to find the volume of both cones then subtract the volume of the smaller container from the larger one area of base for both are equal area of base=area of circle=pi times r^2 radius=diameter/2 diameter=14 radius=14/2 radius=7 pi times 7^2 pi times 49 3.14 times 49=153.86 smaller so 1/3 times 153.86 times 3=smaller cone=153.86 cm^3 larger so 1/3 times 153.86 times 17=871.873 subtract smaller from larger 871.873-153.86=718.013 cm^3 the volume of water needed is 718.01 cm^3

OpenStudy (khalilforthewin):

oh i see what i did wrong

OpenStudy (anonymous):

Oh that's fine. Next question?

OpenStudy (khalilforthewin):

ok